satellite73 (satellite73):

lets try that again can anyone find the range of \[f(x)=x^2-\sqrt{x}\] WITHOUT using calculus?

OpenStudy (anonymous):

can anyone find the range of \[f(x)=x^2-\sqrt{x}\] WITHOUT using calculus?

OpenStudy (anonymous):

i think the domain will be only +ve real numbers ..and zero so ... range should be from zero [0,infi ) i'm not sure but thats what i can make out right now

OpenStudy (anonymous):

\[x=\frac{1}{2}\] \[f(\frac{1}{2})=\frac{1}{4}-\frac{\sqrt{2}}{2}<0\]

OpenStudy (anonymous):

I suppose u could just test 0,1 0,-1 and + -

OpenStudy (anonymous):

i really appreciate your help, satellite73!

OpenStudy (anonymous):

well i don't know how to do this without calc and i don't know what your teacher expects you to do. but the answer you had is wrong.

OpenStudy (anonymous):

satelitte, I have a question, is range from -0.47247 to infinity ?

OpenStudy (amistre64):

the x^2 takes control near infinity so id say its [0,inf) as well

OpenStudy (amistre64):

that came out wrong lol

OpenStudy (amistre64):

ideally id say when sqrt(x) is twice as large as x^2 we could test that point.... sqrt(x) = 2x^2 x = 4x^4 x/x^4 = 4 1/x^3 = 4 x^3 = 1/4 x = cbrt(1/4) .... maybe?

OpenStudy (amistre64):

i spose twice as large superfluous; 5times as large etc amount to the same logic doesnt it

OpenStudy (anonymous):

amistre, check x=0.47247

OpenStudy (anonymous):

I can't find a way to solve it without any calculus, but I can do it using basic notions of limits and extrema, without getting too rigorous or using derivatives: The domain is clearly \[[0, \infty).\]\[f(0) = 0,\ f\left(\frac{1}{2}\right) = \frac{1}{4} - \frac{\sqrt{2}}{2}<0, \ f(4) = 14 > 0\]so there is some interval where the function is decreasing and some interval where the function is increasing. We know that \[x^2 > \sqrt{x}\]if \[x > 1\]and\[x^2 < \sqrt{x}\]if\[0 < x < 1\]so the range of the function is \[(0,\infty)\] on the first interval,\[(1, \infty).\]We know it takes negative values in the second interval and our goal from now on will be to determine how negative those values are. Since\[f(1) = 0,\]the function must start increasing somewhere in\[(0, 1).\]The point where the function starts increasing again will be the minimum (since it's decreasing until that point and then it starts increasing). Let \[1 < \varepsilon < 0\]and\[x < 1\]Suppose \[\varepsilon\]is small enough so \[x + \varepsilon < 1.\]If \[x\]is in the decreasing part of the interval and if \[\varepsilon\]is small enough we must have\[x^2 - \sqrt{x} > (x + \varepsilon)^2 - \sqrt{x + \varepsilon}\](which is the definition of a decreasing function). Squaring what must be squared and tidying up a bit we get\[\varepsilon^2\left(1 - 16 x^3 - 24 x^2 \varepsilon - 12 x \varepsilon^2 + 16 x^4 \varepsilon^2 - 2 \varepsilon^3 + 32 x^3 \varepsilon^3 + 24 x^2 \varepsilon^4 + 8 x \varepsilon^5 + \varepsilon^6\right) > 0.\]Because \[\varepsilon > 0,\]we get\[1 - 16 x^3 + \varepsilon \left(-24 x^2 - 12 x \varepsilon + 16 x^4 \varepsilon - 2 \varepsilon^2 + 32 x^3 \varepsilon^2 + 24 x^2 \varepsilon^3 + 8 x \varepsilon^4 + \varepsilon^5\right) > 0.\]Now, we said\[\varepsilon\]is small. ¿How small? Well, as small as we want it to be. Because x is constant (we aren't changing its value) \[-24 x^2 - 12 x \varepsilon + 16 x^4 \varepsilon - 2 \varepsilon^2 + 32 x^3 \varepsilon^2 + 24 x^2 \varepsilon^3 + 8 x \varepsilon^4 + \varepsilon^5\]is some number that doesn't get too big. Because it's multiplied by \[\varepsilon\]which is a small number, the product is small too (it may be positive or negative, but its absolute value is small). Since this product is arbitrarily small, we can simply discard this number (it doesn't really matter if it's there or not, since it's as small as we want it to be). Then, we get\[1 - 16x^3 > 0\]which finally yields\[x < \frac{1}{\sqrt[3]{16}}\]which is the point where the minimum of the function lies (it's the point where the function stops decreasing and starts increasing). Then, the minimum is\[f\left(\frac{1}{\sqrt[3]{16}}\right) = -\frac{3}{4\sqrt[3]{16^2}}\]so the range of the function is\[\left[-\frac{3}{4\sqrt[3]{16^2}}, \infty\right)\](this is far from a rigorous proof, but I really didn't find a better way without having to use "real" calculus).

OpenStudy (anonymous):

I don't understand that value for f in the last line. If I calculate it, I get about -0.4725 (yours is about -1/8). http://www.wolframalpha.com/input/?i=plot+x^2-sqrt%28x%29

OpenStudy (anonymous):

wow. i didn't mean to create so much work. some one asked this question here (it was in a work sheet) and i said i couldn't do it without calculus but it was a pre-calc course so i thought there was some precalc thing i was missing. i seriously doubt they expected all the work above. totally impressive though

myininaya (myininaya):

\[y=x^2-\sqrt{x}\] \[y-x^2=-\sqrt{x}\] \[(y-x^2)^2=(-\sqrt{x})^2\] \[y^2-2yx^2+x^4=x\] \[y^2-2yx^2+x^4-x=0\] \[x^4-2yx^2+y^2-x=0\] Let u=x^2 \[u^2-2yu+(y^2-x)=0\] \[u=\frac{2y \pm \sqrt{4y^2-4(y^2-x)}}{2}\] we want 4y^2-4(y^2-x)>0 y^2-y^2+x>0 x>0 since the domain of f inverse is x>0 then the range of f is x>0

OpenStudy (anonymous):

Um range is y values...

myininaya (myininaya):

oops y>0

myininaya (myininaya):

but still the graph is lower than 0 so this is not right

OpenStudy (anonymous):

U had right before I think.. "satelitte, I have a question, is range from -0.47247 to infinity ?"

OpenStudy (anonymous):

no but this was something similar to what i had seen a month or so ago. the "solve for x" thing which confused the bejesus out of me. i bet this is right

OpenStudy (anonymous):

who knows? i didn't actually do it. i found the critical point but didn't bother plugging it in to get the decimal

OpenStudy (anonymous):

because that was the calc way and i wanted a different way.

OpenStudy (anonymous):

the critical point is \[\frac{1}{2\sqrt[3]{2}}\] and i didn't feel like trying to evaluate the function there

OpenStudy (anonymous):

and i still don't. but that is the answer, whatever it is \[(\frac{1}{2\sqrt[3]{2}})^2-\sqrt{\frac{1}{2\sqrt[3]{2}}}\]

myininaya (myininaya):

which is approximately -0.4724703927